10 A club consist of ten members. How many different ways can a group of three people be selected to go on an European trip ?
Answer by Alan3354(69443)

A club consist of ten members. How many different ways can a group of three people be selected to go on an European trip ?
-------------------
The 1st selection is 1 of 10, then 1 of 9, then 1 of 8.
10*9*8 = 720.
But, selecting persons ABC is the same as CAB and CBA, etc.
So the 720 has to be divided by 3*2*1 = 6.
720/6 = 120 different possible groups. (Be careful, someone will try to cheat)